OpenStudy (anonymous):

A random sample of 130 students is chosen from a population of 4,500 students. If the mean IQ in the sample is 120 with a standard deviation of 5, what is the 99% confidence interval for the students' mean IQ score?

OpenStudy (anonymous):

118.87−121.13 107.12−132.88 125−135 115−125

OpenStudy (anonymous):

I currently have A as an answer am i right ? @SolomonZelman @cj49 @skullpatrol

OpenStudy (amistre64):

you are going to need to determine a few things, the standard error, and the zscore related to 99%

OpenStudy (amistre64):

\[Z_{99\%}=\frac{x-\bar x}{SE}\] \[Z_{99\%}=\frac{x-\bar x}{\sigma/\sqrt{n}}\]

OpenStudy (anonymous):

am A is correct right

OpenStudy (amistre64):

\[2.576=\frac{CI-120}{5/\sqrt{130}}\] \[120\pm 2.576(5)/\sqrt{130}=CI\]

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=120-2.576%285%29%2Fsqrt%28130%29 118.87 is the lower .. i like it